1a) Calculate the oxidation number for one atom of Mn in the KMnO4 using the reaction: 5Fe2+ + MnO-4 + 8H+\rightarrow 5Fe3+ +Mn2+ + H2O

1b) Considering the change in oxidation number of Mn in the product, how many electrons are gained or lost by Mn in KMnO4 when the reaction is balanced?
Answer #1

Oxidation number of Mn in KMnO4;

1 + x + 2*(-4) = 0

x = +7

OR

Reactant side charge = product side charge

5*(2) + x + 4*(-2) + 8 = 17

10 + x - 8 + 8 = 17

x = +7

In product side, Mn2+

Oxidation number of Mn is +2

Gain of e- = 7 - 2 = 5 ...Answer